OpenStudy (anonymous):

how to calculate integration of 1/(x+2) ki power 2

OpenStudy (mimi_x3):

\[ \int \frac{1}{(x+2)^2}dx\]???

OpenStudy (mimi_x3):

if yes, let \(u = x+2\)

OpenStudy (mimi_x3):

\[\int \frac{1}{u^2}du => \int u^{-2}du\]

OpenStudy (mimi_x3):

\[=>\frac{u^{-1}}{-1}+c\] Then sub back the \(u\)

OpenStudy (anonymous):

if it is 6/(x+2) ki power 2 then

OpenStudy (mimi_x3):

what is \(ki\)

OpenStudy (anonymous):

i wrote the question

OpenStudy (mimi_x3):

hmm..i dont understand; what is \(ki\) first??

OpenStudy (anonymous):

6/(x+2) ki power 2

OpenStudy (anonymous):

ka integration

OpenStudy (anonymous):

6/(x+2) ki power 2 ka integration, pls help

OpenStudy (mimi_x3):

so \[\int \frac{6}{x+2}du\] u = \(x+2\) => \[\frac{du}{dx} =2\]

hartnn (hartnn):

you mean [6/(x+2) ]^2 dx or 6 / (x+2)^2 dx ?

OpenStudy (anonymous):

upar wala

hartnn (hartnn):

means the upper one

hartnn (hartnn):

english me bol na

OpenStudy (anonymous):

yes

hartnn (hartnn):

its same as your original Q just 6 is multiplied -_-

OpenStudy (anonymous):

but the answer does not match by that way why? and the answer is 3x/x+2

OpenStudy (anonymous):

guru hoza shuru jadu ki chhadi ghumao aur answer batao

hartnn (hartnn):

integral of 6 / (x+2)^2 dx is -6/(x+2). and its correct.

hartnn (hartnn):

-6/(x+2) +c

OpenStudy (anonymous):

that means the answer behind the book is wrong

hartnn (hartnn):

must be